A local ice hockey team has asked you to design an
Answer #1

Part-A : If the rod makes one revolution every 0.736 s after the puck is caught, then the puck's speed just before it hit the rod which is given as -

\omegaf = (1 x 2\pi) / (0.736 s) = 8.53 rad/s

using conservation of momentum, we have

mr v0,r + mp v0,p = (mr + mp) vf

where, v0,r = initial speed of rod = 0 m/s

we know that, angular velocity is given by -

\omegaf = vf / r \Leftrightarrow (8.53 rad/s) = vf / (2 m)

vf = 17.06 m/s

then, we get

(1.3 kg) (0 m/s) + (0.163 kg) v0,p = [(1.3 kg) + (0.163 kg)] (17.06 m/s)

(0.163 kg) v0,p = (24.9 kg.m/s)

v0,p = 152.7 m/s